A atoms at corners B atoms at face centre find formula of compound
 

Grade:12

2 Answers

Yash
4 years ago
AB3
Number of A atoms from 8 corners  is 1
Number of B atoms from 6 faces is 3
Therefore formula of compound is AB3
Khimraj
4 years ago
 
Number of A atoms from 8 corners  = 8*1/8 =1
Number of B atoms from 6 faces is = 6*1/2 = 3
Therefore formula of compound is AB3